The difference between two numbers is five. The larger number is equal to one more than double the smaller number. Write two equations then solve them to find the two numbers.

Answer by CubeyThePenguin(3113)

x - y = 5
x = 2y + 1

Substitute the second equation into the first equation.

(2y + 1) - y = 5
y + 1 = 5
y = 4

The integers are x = 9 and y = 4.